The equilibrium constant, Kc, for the following reaction is 1.80x102 at 698 K. Calculate the equilibrium concentrations of reactant and products when 0.292 moles of HI are introduced into a 1.00 L vessel at 698 K HI] 121
Answer #1

[HI]= = 0.292 _ 0-292M 2 HI H I Elo.292 - 2x [H] [I] kr {HI} 1.8x10 O = CO-292-24) 1.8x10 = 1.085264+ 4x²_1-1684 1.88102 (0.01024x1020-0783518 -2.1024x10 +0.0783518 Xa- X = 1-856 1-856 x = 0.030 888 M ✓ x=tive o [#3 = [I= 30-030888 M [HI] = 0.292-24